# Restivo Salemi property for $\alpha$-power free languages with $\alpha\geq 5$ and $k\geq 3$ letters

### v1

@@ 4 Transition words
To get a contradiction, suppose that there are $g\in\operatorname{Prf}(\widetilde{w}_{2}xt_{2})$, $y\in\Sigma_{k}$, $r\in\Sigma_{k}^{+}$, and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(s_{1}\widetilde{w}_{1}hpgy)$, $gy\in\operatorname{Prf}(\widetilde{w}_{2}xt_{2})$, and $s_{1}\widetilde{w}_{1}hpg\in L_{k,\alpha}^{\mathbb{N},L}$. Since $p\widetilde{w}_{2}xt_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $x\in F(w_{2})$ we have that $|r^{\beta}|>|pgy|$ and $x\in F(r)$.
@@
Since $p\widehat{w}_{2}x\in F(r^{\beta})$, $x\not\in F(p\widehat{w}_{2})$, and $|p|>|\widetilde{w}_{2}|$ we have that
@@
Recall $x\not\in F(s_{1})\cup F(s_{2})\cup F(t_{1})\cup F(t_{2})$. Then it follows from (1) that
$\operatorname{occur}(s_{1}\widetilde{w}_{2}hpw_{2}t_{2},p\widehat{w}_{2}x)\leq 2\mbox{.}$ (3)
From (2), (3), and $\beta>\alpha\geq 4$ we conclude that there are no such $r,g,\beta$ and hence $s_{1}x\widetilde{w}_{1}xhpx\widetilde{w}_{2}xt_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
@@
The main theorem of the article says that if $w_{1},w_{2}$ are finite nonempty factors of bi-infinite $\alpha$-power free words then there is a bi-infinite $\alpha$-power free word $v$ and a finite word $w$ such that $w_{1}ww_{2}$ is a factor of $v$.
@@
If $v_{1},v_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $w_{1}\in F(v_{1})\setminus\{\epsilon\}$, and $w_{2}\in F(v_{2})\setminus\{\epsilon\}$ then there are $w\in\Sigma_{k}^{*}$ and $v\in L_{k,\alpha}^{\mathbb{Z}}$ such that $w_{1}ww_{2}\in F(v)$.
@@
If $F(v_{1})\cap F(v_{2})=\emptyset$ then let $v_{11}$, $v_{12}$, $v_{21}$, and $v_{22}$ be such that $v_{1}=v_{11}v_{12}$, $v_{2}=v_{21}v_{22}$, $w_{1}\in F(v_{11})$, and $w_{2}\in F(v_{22})$. Obviously $v_{11}{v_{22}}$ is $\alpha$-power free since $F(v_{11})\cap F(v_{22})=\emptyset$.
Thus suppose that $F(v_{1})\cap F(v_{2})\not=\emptyset$ and let $x\in F(v_{1})\cap F(v_{2})\cap\Sigma_{k}$. Without loss of generality we may suppose that $x\in F(w_{1})\cap F(w_{2})$.
Lemma 4.2 asserts that there are $s_{1}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma_{1},\eta_{1}\in\Sigma_{k}^{*}$ such that $(s_{1},\sigma_{1},w_{1},\eta_{1},x)\in\Delta$. Lemma 4.2 asserts also that there are $t_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\sigma_{2},\eta_{2}\in\Sigma_{k}^{*}$ such that $(t_{2}^{R},\sigma_{2}^{R},w_{2}^{R},\eta_{2}^{R},x)\in\Delta$.
Let $t_{1}$ be a right infinite $\alpha$-power free word on the alphabet $\Sigma_{k}\setminus\{x\}$. Since $\alpha\geq 4$, Lemma 2.3 asserts that $t$ exists. Let $s_{2}\in\Phi(t)$; Lemma 2.4 implies that $\Phi(t_{1})\not=\emptyset$. Clearly $s_{2}$ is $\alpha$-power free and $x\not\in F(s_{2})$.
Theorem 4.1 implies that are $\overline{\eta}_{1}\in\operatorname{Prf}(\eta_{1})$ and $\overline{\eta}_{2}\in\operatorname{Suf}(\eta_{2})$ such that $s_{1}\sigma_{1}w_{1}\overline{\eta}_{1}xt_{1},t_{2}^{R}\sigma_{2}^{R}w_{2}^{R}\overline{\eta}_{2}^{R}xs_{2}^{R}\in L_{k,\alpha}^{\mathbb{Z}}$. Since the reverse operation preserves the power-freeness we have also that $s_{2}x\overline{\eta}_{2}w_{2}\sigma_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
Let $\widetilde{w}_{1}=\sigma_{1}w_{1}\overline{\eta}_{1}x$ and let $\widetilde{w}_{2}=x\overline{\eta}_{2}w_{2}\sigma_{2}$. Then Proposition 4.3 implies that there is $w\in\Sigma_{k}^{*}$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$. The theorem follows.
@@
$\overline{w}_{1}$ $\overline{w}_{2}$
$s_{1}$ $\sigma_{1}$ $w_{1}$ $\overline{\eta}_{1}$ $x$ $h$ $p$ $x$ $\overline{\eta}_{2}$ $w_{2}$ $\sigma_{2}$ $t_{2}$
Table 1: The structure of the transition word $v$.
This completes the proof. ∎

### v4

@@ 3 Bi-infinite $\alpha$-power words
To get a contradiction, suppose that there are $g\in\operatorname{Prf}(\widetilde{w}_{2}t_{2})$, $y\in\Sigma_{k}$, $r\in\Sigma_{k}^{+}$, and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(s_{1}\widetilde{w}_{1}hpgy)$, $gy\in\operatorname{Prf}(\widetilde{w}_{2}t_{2})$, and $s_{1}\widetilde{w}_{1}hpg\in L_{k,\alpha}^{\mathbb{N},L}$. Since $p\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $x\in\operatorname{Prf}(\widetilde{w}_{2})$ we have that $|r^{\beta}|>|pgy|$ and $x\in F(r)$.
@@
Since $p\widehat{w}_{2}x\in F(r^{\beta})$ and $x\not\in F(p\widehat{w}_{2})$, we have that
@@
Recall $x\not\in F(s_{1})\cup F(hp)\cup F(t_{2})$. Then it follows from (4) that
$\operatorname{occur}(s_{1}\widetilde{w}_{1}hp\widetilde{w}_{2}t_{2},p\widehat{w}_{2}x)\leq 2\mbox{.}$ (6)
From (5), (6), and $\beta>\alpha\geq 5$ we conclude that there are no such $r,g,\beta$ and hence $s_{1}\widetilde{w}_{1}hp\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
@@
The main theorem of the article says that if $w_{3},w_{4}$ are finite non-empty factors of bi-infinite $\alpha$-power-free words then there is a bi-infinite $\alpha$-power-free word $v$ and a finite word $w_{0}$ such that $w_{3}w_{0}w_{4}$ is a factor of $v$.
@@
If $v_{1},v_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $w_{3}\in F(v_{1})\setminus\{\epsilon\}$, and $w_{4}\in F(v_{2})\setminus\{\epsilon\}$ then there are $w_{0}\in\Sigma_{k}^{*}$ and $v\in L_{k,\alpha}^{\mathbb{Z}}$ such that $w_{3}w_{0}w_{4}\in F(v)$.
@@
If $F(v_{1})\cap F(v_{2})=\emptyset$ then let $v_{11}$, $v_{12}$, $v_{21}$, and $v_{22}$ be such that $v_{1}=v_{11}v_{12}$, $v_{2}=v_{21}v_{22}$, $w_{3}\in F(v_{11})$, and $w_{4}\in F(v_{22})$. It is clear $v_{11}{v_{22}}$ is $\alpha$-power-free since $F(v_{11})\cap F(v_{22})=\emptyset$. The theorem follows.
Thus suppose that $F(v_{1})\cap F(v_{2})\not=\emptyset$ and let $x\in F(v_{1})\cap F(v_{2})\cap\Sigma_{k}$. Let $w_{1}\in F(v_{1})$ and $w_{2}\in F(v_{2})$ be such that $x\in F(w_{1})\cap F(w_{2})$, $w_{3}\in F(w_{1})$, and $w_{4}\in F(w_{2})$. It is clear that such $w_{1},w_{2}$ exist.
Lemma 3.7 asserts that there are $s_{1}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma_{1},\eta_{1}\in\Sigma_{k}^{*}$ such that $(s_{1},\sigma_{1},w_{1},\eta_{1},x,\epsilon)\in\Delta$. Lemma 3.7 asserts also that there are $t_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\sigma_{2},\eta_{2}\in\Sigma_{k}^{*}$ such that $(t_{2}^{R},\sigma_{2}^{R},w_{2}^{R},\eta_{2}^{R},x,\epsilon)\in\Delta$.
Let $t_{1}\in L_{k-1,\alpha}^{\mathbb{N},R}$ be a right infinite $\alpha$-power-free word on the alphabet $\Sigma_{k}\setminus\{x\}$. Since $\alpha\geq 5$, Lemma 2.4 asserts that $t_{1}$ exists. Let $s_{2}\in\Phi(t_{1})$; Lemma 2.5 implies that $\Phi(t_{1})\not=\emptyset$ and consequently $s_{2}$ exists. Clearly $s_{2}$ is $\alpha$-power-free and $x\not\in F(s_{2})$.
Theorem 3.4 implies that there are $\widehat{\eta}_{1}\in\operatorname{Prf}(\eta_{1})$ and $\widehat{\eta}_{2}\in\operatorname{Suf}(\eta_{2})$ such that $s_{1}\sigma_{1}w_{1}\widehat{\eta}_{1}xt_{1},t_{2}^{R}\sigma_{2}^{R}w_{2}^{R}\widehat{\eta}_{2}^{R}xs_{2}^{R}\in L_{k,\alpha}^{\mathbb{Z}}$. Since the reverse operation preserves the power-freeness we have also that $s_{2}x\widehat{\eta}_{2}w_{2}\sigma_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
Let $\widetilde{w}_{1}=\sigma_{1}w_{1}\widehat{\eta}_{1}x$ and let $\widetilde{w}_{2}=x\widehat{\eta}_{2}w_{2}\sigma_{2}$. Then Proposition 3.8 implies that there is $w\in\Sigma_{k}^{*}$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
@@
$\widetilde{w}_{1}$ $\widetilde{w}_{2}$
$s_{1}$ $\sigma_{1}$ $w_{1}$ $\widehat{\eta}_{1}$ $x$ $h$ $p$ $x$ $\widehat{\eta}_{2}$ $w_{2}$ $\sigma_{2}$ $t_{2}$
Table 1: The structure of the word $v$.
Since $w_{3}\in F(w_{1})$, $w_{4}\in F(w_{2})$, $w_{1}\in F(\widetilde{w}_{1})$, and $w_{2}\in F(\widetilde{w}_{2})$, the theorem follows. This completes the proof. ∎
Theorem 3.9 has the following obvious corollary.
Corollary 3.10.
Conjecture 1.3 holds for the languages $L_{k,\alpha}$, where $\alpha\geq 5$ and $k\geq 3$.